OpenStudy (anonymous):

verify the identity cot^2x + csc^2x = 2csc^2x - 1 please help me and i dont just want the work i want to understand how to do it.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

rewrite everything in terms of sine and cosine, then do the arithmetic on the left, may work

OpenStudy (anonymous):

can you walk me through it @misty1212

OpenStudy (misty1212):

lets call cosine \(a\) and call sine \(b\) then the left hand side is \[\frac{a^2}{b^2}+\frac{1}{b^2}\]

OpenStudy (misty1212):

add them up and you get \[\frac{a^2+1}{b^2}\]

OpenStudy (anonymous):

where are you getting cosine and sine

OpenStudy (anonymous):

nvm this is easier than all that

OpenStudy (anonymous):

\[\cot^2(x)=\csc^2(x)-1\] is all we can prove that if you like

OpenStudy (anonymous):

start with \[\sin^2(x)+\cos^2(x)=1\] then divide all by \(\sin^2(x)\) to ge t \[1+\cot^2(x)=\csc^2(x)\]

OpenStudy (anonymous):

that means \[\cot^2(x)=\csc^2(x)-1\]

OpenStudy (anonymous):

plug that in and you get exactly what you want

OpenStudy (anonymous):

\[\cot^2(x)+\csc^2(x)=\csc^2(x)+\csc^2(x)-1=2\csc^2(x)-1\] done

OpenStudy (anonymous):

hey sorry @satellite73

OpenStudy (owlcoffee):

\[\cot^2x+\csc^2=2\csc^2x-1\] In order to prove this equality, I will just focus on the left side: \[\cot^2x+\csc^2\] And I will try to make them simple by using the equality: \[\cot^2x=\csc^2x-1\] So, then: \[(\csc^2x-1)+\csc^2x\] And now it's just a matter of simplying by operating the similar factors, I'll call csc^2x=u just to make it simpler for you: \[u-1+u\] And then, operating: \[2u-1\] And, disarming the change of variable: \[2\csc^2x-1\] And that proves it.